A car drives on a road which begins flat, then goes up and over the top of a hill and then goes down into a dip and back up to level ground.

12)  If the car is moving at 30 m/s at the bottom of the dip, its mass is 800 kg, and the radius of curvature for the dip is 20 m, what is the normal force acting on the car in the dip.

10459 N

19344 N

28625 N

37190 N

43840 N

13)  What is the acceleration of the car at the bottom of the dip?

0 m/s2

9.8 m/s2

35.2 m/s2

45.0 m/s2

54.8 m/s2
Answer #1

m = 800 kg

v = 30 m/s

r = 20 m

12) F = mg + mv^2/r = 800*9.8 + (800*900)/20 = 43840 N

13)  a = F/m = 43840/800 = 54.8 m/s2